# Strongly Coupled Sectors in Inflation: Gapless Theories and Unparticles

### v3

@@ 1 Introduction
Under the assumption that the universe went through a phase of cosmic inflation, there are many ways of embedding the scalar perturbations in a more complete theory. The effective theory of inflation [Cheung:2007st, Baumann:2009ds, Lee:2016vti] is similar to that of the non-linear sigma model. One way of ultraviolet (UV) completing this theory is to consider a linear sigma model in a broken symmetry phase—this is what happens within the Standard Model for the Higgs. The analogous mechanism in early-universe cosmology falls under the umbrella of Single Field Inflation. Of course, we can then consider more sophisticated scenarios, where the inflaton is weakly coupled to other particles, like in the electroweak sector of the Standard Model. In recent years, this class of scenarios [^1] In its simplest incarnation, models of spectators weakly coupled to the inflaton are called “quasi-single field inflation” [Chen:2009we, Chen:2009zp, Chen:2010xka, Baumann:2011nk, Assassi:2012zq, Noumi:2012vr]. It was then emphasized that the states in the UV completion can be determined by precise shapes of non-Gaussianities, thus making the analogy with collider physics quantitative—what is now referred to as “cosmological collider physics” [Arkani-Hamed:2015bza]. have been studied in much detail [Maldacena:2002vr, Chen:2010xka, Arkani-Hamed:2015bza, Arkani-Hamed:2018kmz, Baumann:2019oyu, Baumann:2020dch, Aoki:2023tjm]. In this direction, the frontier of research is to do precision calculations, either at tree-level and high-point functions, or at loop level. The resulting shapes of primordial non-Gaussianity, so-called cosmological correlators, are now well-understood, from their analytic structure to phenomenological consequences.
The alternative way of UV completing the non-linear sigma model is to think of the scalar fluctuation as a composite particle, made out of yet more elementary constituents. This is what happens in Quantum Chromodynamics (QCD) and in models of compositeness for the Higgs [Weinberg:1975gm, Susskind:1978ms, Kaplan:1983fs, Kaplan:1983sm, Dugan:1984hq, Barbieri:2007bh, Bellazzini:2014yua, Panico:2015jxa, Goertz:2018dyw], etc. Such a scenario has been way less explored in this direction; for a few recent examples in this direction see e.g. [Kumar:2018jxz, Hubisz:2024xnj]. We will consider a simpler situation. We couple the curvature perturbations weakly to a separate sector, and enlarge the non-linear sigma model to accommodate this coupling. This separate sector is strongly coupled. The idea is that this is a first step to understand general features of strongly coupled dynamics in inflation. Of course it would be interesting to have benchmark examples where background evolution and the scalar curvature perturbation come from compositeness, but we will not attempt to do that here.
@@
  • $\text{M}_{\text{Gap}}\ll H$: If the theory has a very small mass gap, we could approximate it to be a four-dimensional conformal field theory (CFT4). It will have operators with nontrivial anomalous dimensions—so-called “unparticles,” as well as of course a stress tensor, and possibly conserved currents more generally. This scenario will be the main focus of our paper. It has been studied in [Green:2013rd] (see also the recent [Chakraborty:2025myb]), where many general features were described, mostly on squeezed and collapsed limits. We will review some of them, and compute the full shapes of the corresponding cosmological correlator using various techniques.
  • $\text{M}_{\text{Gap}}\sim H$. Inflation is a good particle detector for masses below, at and slightly above the Hubble scale.  [^2] In scenarios with a chemical potential [Flauger:2016idt, Bodas:2020yho], one can probe energy scales up to the strong coupling scale for the pions, which is $\sim 60H$ in slow-roll inflation. In this case, many bound states of the strongly coupled sector, with masses of order Hubble, will be produced and propagate for a few Hubble volumes before their power gets diluted. They will interfere in an interesting way, potentially making the resulting imprint on cosmological correlators to be novel, compared to the usual quasi-single field inflation.
@@
  • Sections 4.2 and 4.3 show that when $\Delta\geq 3/2$, both trispectra and bispectra in the soft limit are degenerate with the equilateral non-Gaussianity. This fact requires us to look into the full shape, instead of the collapsed/squeezed limit, to search for unparticles. For $\Delta<3/2$ the squeezed limits are degenerate with exchange of particles in the complementary series. Once again, only a comparison of full shapes can conclusively break this degeneracy.
@@ 2 Setup
In this section, we briefly discuss a setup that accommodates strongly coupled spectator fields in inflation, of various mass gaps. If the theory is strongly coupled and has a very large number of degrees of freedom, then we know that its supersymmetric versions have a weakly coupled dual description that is gravitational and holographic [Hawking:1982dh, Banados:1997df, Maldacena:1997re, Witten:1998qj, Banados:1998dc, Witten:1998zw, Buchel:2002wf, Buchel:2002kj, Aharony:2002cx, Balasubramanian:2002am, Cai:2002mr, Hertog:2004rz, Ross:2004cb, Alishahiha:2004md, Hertog:2005hu, Balasubramanian:2005bg, Hirayama:2006jn, He:2007ji, Cacciapaglia:2007jq, Cacciapaglia:2008ns, Hutasoit:2009xy, Marolf:2010tg, Maldacena:2010un, Chu:2016pea, Kumar:2018jxz, Csaki:2018kxb, Megias:2019vdb, Megias:2021arn, Penin:2021sry, Mishra:2022fic]. In that case we can also describe its holographic dual. We make a few comments on these holographic duals, mostly to illustrate that the resulting cosmological correlators will have new features. Then we go back to our main focus, that of gapless sectors.
@@ 2.1 Unparticle Physics in Inflation
The infrared (IR) phases of quantum field theories (QFT) can be generally classified into three possibilities: a theory with a mass gap, a theory with massless particles in the IR or a scale-invariant theory with a continuous spectrum [Rychkov:2016iqz]. In particle phenomenology, a hypothetical scale-invariant field is called unparticle, introduced in [Banks:1981nn, Georgi:2007ek, Georgi:2007si, Georgi:2008pq, Georgi:2009xq, Grinstein:2008qk, Stephanov:2007ry]. It is reasonable to consider the unparticles to be conformal fields since we want to capture general features of this scale-invariant sector. Some examples of Beyond the Standard Model (BSM) physics that are conformal or approximately conformal include the second Randall-Sundrum model [Randall:1999vf], the “hidden-valley” scenario [Strassler:2006im], and the conformal window in QCD [Gardi:1998ch]—all of them can be regarded as “unparticle physics”.
It is possible that in the primordial universe, the UV theory contains some fields that become non-trivially scale-invariant as we probe the Hubble scale during inflation. Typically they will become strongly coupled and the operators in the spectrum are composites of the UV fields [Banks:1981nn, Sannino:2008nv]. This non-trivial scaling invariant sector can possibly be discovered experimentally in some missing energy distributions, for example in [Fox:2011pm, CMS:2014jvv, Sannino:2009za, Cheung:2007zza, Cheung:2007ap], and in cosmological observations, discussed in [Kikuchi:2007az, Collins:2008ny, McDonald:2008uh, Green:2013rd, Baumgart:2021ptt].
Let’s review how to couple operators to the metric within the effective field theory (EFT) of inflation, e.g. [Cheung:2007st, Lee:2016vti]. Inflation introduces a natural foliation of spacetime with a clock, so the time-components of the metric $g^{00}$ become acceptable operators on their own.  [^3] Within single-field slow-roll inflation, one can see this by simply evaluating the kinetic term of the inflaton on its background value $\bar{\phi}$, $(\nabla\phi)^{2}\to g^{00}\dot{\bar{\phi}}^{2}$. In this unitary gauge, for a general action in an Friedmann–Robertson–Walker (FRW) universe constructed from products of any four-dimensional covariant tensors with free upper 0 indices, in the leading order, we can write out
@@
Neglecting the linear term above certain energy [Cheung:2007st], now we construct the interactions when we couple the unparticles to the Goldstone $\pi$. In this work, we will be most interested in the interactions between the Goldstone and the scalar unparticles. We will also consider the couplings to the unparticle current and the stress tensor of unparticles. Since the unparticles are scaling invariant, the FRW curvature will not affect the conservation of the current, while it is natural to consider the traceless part of the stress tensor.
@@
where $a(t)$ is the FRW scaling factor and we used the constraint $\nabla^{\mu}J_{\mu}=0$ at the background level, to replace $\dot{\pi}J_{0}$ by $J_{i}\partial_{i}\pi$. For spin-2 traceless stress tensor field $T_{\mu\nu}\equiv\mathcal{O}^{(2)}_{\mu\nu,\Delta=4}$, similar to the above discussion, the leading order interactions between $T_{\mu\nu}$ and the Goldstone $\pi$ will be
@@
where $f_{\pi}^{4}$ is the symmetry breaking scale [Baumann:2011su] and we show the amplitude of scalar fluctuations that best fits the Planck satellite observations [Planck:2018vyg].
@@ 2.2 Strongly Coupled Sectors with Holographic Duals
It is of course difficult to treat strongly coupled field theories, as there are no general tools to study them. Besides going to the lattice, a very useful alternative is to use a weakly coupled dual description, if available. For some theories, the dual description is gravitational. We are considering a case in which we decouple gravity from the scalar dynamics during inflation, and consider a strongly coupled field theory in a rigid background. In that case, we can approximate the inflationary geometry to a fixed de Sitter background, and describe the holographic duals of strongly coupled quantum field theories in de Sitter space, see e.g. the discussion in [Marolf:2010tg, Maldacena:2012xp, Hubisz:2024xnj].
@@
Now we consider mass deformations of the UV fixed point. The resulting gravity duals will depend on the dominant phase of the theory in the large $N$ limit [Hawking:1982dh, Witten:1998zw]. We will content ourselves with describing the gravitational duals that correspond to each phase, though at finite $N$ they will both be present and mixed by thermal fluctuations. The scenarios are similar to those described in [Randall:1999ee, Randall:1999vf, Maldacena:2010un, Maldacena:2012xp]. A useful setup is that of a five-dimensional CFT compactified on $\text{dS}_{4}\times S^{1}$. There are two saddles. The first one is the Schwarzschild geometry, which when Wick rotated gives a bubble of nothing spacetime [Witten:1998zw, Balasubramanian:2002am, Maldacena:2012xp], or a spacetime with an end of the world brane. This is the gravity dual of a gapped theory with mass much bigger than the Hubble scale of the boundary. As the mass gap is rather big, the states are inefficiently produced (being Boltzmann suppressed), so we expect the cosmological correlator in this phase to be given by tree-level exchange of the lightest state of the spectrum—which, being heavy, will be well-approximated by the equilateral shape of non-Gaussianity.
@@
Finally, when the mass deformation is not too large, there is a phase of the theory with a gap around the Hubble scale. The gravity dual has a holographic renormalization group (RG) flow that can be continued past the location of the putative IR brane, effectively “frying it.” [^4] Thanks to Raman Sundrum for explaining the appearance of the horizon through “brane frying.” In that case, the dual develops a cosmology behind this brane, a crunching cosmology with open FRW slices—see the discussion in [Maldacena:2012xp]. A simple model of this gravity dual is $\text{AdS}_{6}$ with an identification. The identification makes the compactified circle shrink to zero size at the crunch.
@@
From [Arkani-Hamed:2018kmz, Baumann:2019oyu], we know that we can relate the correlation functions of massless inflatons $\phi$ to the correlation functions of conformally coupled scalars $\varphi$ through weight-shifting operators. Thus we consider an effective action of a conformally coupled $\varphi$ coupled to an unparticle field of dimension $\Delta$:
@@
In inflationary cosmology, since we are computing the equal-time correlator on the late-time boundary, we go to spatial momentum space. From [Grinstein:2008qk, Bautista:2019qxj], we know that in $d$-dimensional Euclidean space,
@@
after normalizing this expression by $(2\pi)^{2}$. If we take $\Delta=1$, (3.11) will reduce to the well-known propagator of conformally coupled scalars [Arkani-Hamed:2015bza]:
@@
In the “in-in” formalism (also called Schwinger-Keldysh formalism), the four-point correlation function with two-site tree-level exchange in the $s$-channel can always be divided into four sectors [Schwinger:1960qe, Keldysh:1964ud, Jordan:1986ug, Calzetta:1986ey, Maldacena:2002vr, Chen:2010xka, Arkani-Hamed:2015bza]:
@@ 3.1 Mellin-Barnes Integration
To make the notation consistent with [Arkani-Hamed:2018kmz], we map $X_{1}\mapsto 1/u$ and $X_{2}\mapsto 1/v$. Now the physical regions of the kinematic variables are $u\in(0,1]$, $v\in(0,1]$. Based on the derivation explained in detail in Appendix B, using Mellin-Barnes integration, we obtain the analytic form of the de Sitter four-point function, given by
@@
When $\Delta$ is an integer, there will be UV-divergences in (3.22). Physically, these divergences occurring at integer values of $\Delta$ can be interpreted as the divergences in $n$-loop “banana” diagrams involving multiple propagators. This is explained in detail in [Westerdijk:2025ywh]; the essential idea is that in free field theory one can build integer-valued unparticle operators from taking operator products of free fields. This procedure requires proper renormalization, which is reflected in the divergences of the unparticle exchange diagram. Moreover, [Bautista:2019qxj] showed that when $\Delta-2\in\mathds{N}^{0}$, we need to regularize the 2-point function, because the Euclidean 2-point function has no Fourier transform and requires regularization and renormalization, which leads to anomalies. The UV-divergences can be absorbed in local counterterms. For four-point functions, the counterterms we are able to add to the Lagrangian will be the contact interaction $\lambda\varphi^{4}$ and its higher-derivatives. As we will illustrate with the following examples, these counterterms are able to absorb the UV divergences. From [Arkani-Hamed:2018kmz], we know that contact interactions from irrelevant operators have shapes that can be calculated by repeated application of $\Delta_{u}$ on the simplest correlator coming from a $\varphi^{4}$ self-interaction, given by $uv/(u+v)$, with
@@
To perform the renormalization for integer $\Delta$’s, one has to expand the Appell $F_{1}$ functions. With the help of the package Diogenes developed in [Bezuglov:2023owj], we can find $\epsilon$-expansions for Appell $F_{1}$ functions expanded near integer values. The $\Delta=1$ case is well-known [Arkani-Hamed:2015bza, Arkani-Hamed:2018kmz], it is given by

### v4

@@ 1 Introduction
Under the assumption that the universe went through a phase of cosmic inflation, there are many ways of embedding the scalar perturbations in a more complete theory. The effective theory of inflation [1, 2, 3] is similar to that of the non-linear sigma model. One way of ultraviolet (UV) completing this theory is to consider a linear sigma model in a broken symmetry phase—this is what happens within the Standard Model for the Higgs. The analogous mechanism in early-universe cosmology falls under the umbrella of Single Field Inflation. Of course, we can then consider more sophisticated scenarios, where the inflaton is weakly coupled to other particles, like in the electroweak sector of the Standard Model. In recent years, this class of scenarios [^1] In its simplest incarnation, models of spectators weakly coupled to the inflaton are called “quasi-single field inflation” [4, 5, 6, 7, 8, 9]. It was then emphasized that the states in the UV completion can be determined by precise shapes of non-Gaussianities, thus making the analogy with collider physics quantitative—what is now referred to as “cosmological collider physics” [10]. have been studied in much detail [11, 6, 10, 12, 13, 14, 15]. In this direction, the frontier of research is to do precision calculations, either at tree-level and high-point functions, or at loop level. The resulting shapes of primordial non-Gaussianity, so-called cosmological correlators, are now well-understood, from their analytic structure to phenomenological consequences.
The alternative way of UV completing the non-linear sigma model is to think of the scalar fluctuation as a composite particle, made out of yet more elementary constituents. This is what happens in Quantum Chromodynamics (QCD) and in models of compositeness for the Higgs [16, 17, 18, 19, 20, 21, 22, 23, 24], etc. Such a scenario has been way less explored in this direction; for a few recent examples in this direction see e.g. [25, 26]. We will consider a simpler situation. We couple the curvature perturbations weakly to a separate sector, and enlarge the non-linear sigma model to accommodate this coupling. This separate sector is strongly coupled. The idea is that this is a first step to understand general features of strongly coupled dynamics in inflation. Of course it would be interesting to have benchmark examples where background evolution and the scalar curvature perturbation come from compositeness, but we will not attempt to do that here.
@@
  • $\text{M}_{\text{Gap}}\ll H$: If the theory has a very small mass gap, we could approximate it to be a four-dimensional conformal field theory (CFT4). It will have operators with nontrivial anomalous dimensions—so-called “unparticles,” as well as of course a stress tensor, and possibly conserved currents more generally. This scenario will be the main focus of our paper. It has been studied in [27] (see also the recent [28]), where many general features were described, mostly on squeezed and collapsed limits. We will review some of them, and compute the full shapes of the corresponding cosmological correlator using various techniques.
  • $\text{M}_{\text{Gap}}\sim H$. Inflation is a good particle detector for masses below, at and slightly above the Hubble scale.  [^2] In scenarios with a chemical potential [29, 30], one can probe energy scales up to the strong coupling scale for the pions, which is $\sim 60H$ in slow-roll inflation. In this case, many bound states of the strongly coupled sector, with masses of order Hubble, will be produced and propagate for a few Hubble volumes before their power gets diluted. They will interfere in an interesting way, potentially making the resulting imprint on cosmological correlators to be novel, compared to the usual quasi-single field inflation.
@@
  • Sections 4.2 and 4.3 show that, when $\Delta\geq 3/2$ for trispectra and $\Delta\geq 2$ for bispectra, the soft limit behaviors are degenerate with the equilateral non-Gaussianity. This fact requires us to look into the full shape, instead of the collapsed/squeezed limit, to search for unparticles. For $\Delta<2$, the squeezed limit degenerates with the exchange of particles in complementary series. Once again, only a comparison of full shapes can conclusively break this degeneracy.
@@ 2 Setup
In this section, we briefly discuss a setup that accommodates strongly coupled spectator fields in inflation, of various mass gaps. If the theory is strongly coupled and has a very large number of degrees of freedom, then we know that its supersymmetric versions have a weakly coupled dual description that is gravitational and holographic [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 25, 55, 56, 57, 58, 59]. In that case we can also describe its holographic dual. We make a few comments on these holographic duals, mostly to illustrate that the resulting cosmological correlators will have new features. Then we go back to our main focus, that of gapless sectors.
@@ 2.1 Unparticle Physics in Inflation
The infrared (IR) phases of quantum field theories (QFT) can be generally classified into three possibilities: a theory with a mass gap, a theory with massless particles in the IR or a scale-invariant theory with a continuous spectrum [60]. In particle phenomenology, a hypothetical scale-invariant field is called unparticle, introduced in [61, 62, 63, 64, 65, 66, 67]. It is reasonable to consider the unparticles to be conformal fields since we want to capture general features of this scale-invariant sector. Some examples of Beyond the Standard Model (BSM) physics that are conformal or approximately conformal include the second Randall-Sundrum model [68], the “hidden-valley” scenario [69], and the conformal window in QCD [70]—all of them can be regarded as “unparticle physics”.
It is possible that in the primordial universe, the UV theory contains some fields that become non-trivially scale-invariant as we probe the Hubble scale during inflation. Typically they will become strongly coupled and the operators in the spectrum are composites of the UV fields [61, 71]. This non-trivial scaling invariant sector can possibly be discovered experimentally in some missing energy distributions, for example in [72, 73, 74, 75, 76], and in cosmological observations, discussed in [77, 78, 79, 27, 80].
Let’s review how to couple operators to the metric within the effective field theory (EFT) of inflation, e.g. [1, 3]. Inflation introduces a natural foliation of spacetime with a clock, so the time-components of the metric $g^{00}$ become acceptable operators on their own.  [^3] Within single-field slow-roll inflation, one can see this by simply evaluating the kinetic term of the inflaton on its background value $\bar{\phi}$, $(\nabla\phi)^{2}\to g^{00}\dot{\bar{\phi}}^{2}$. In this unitary gauge, for a general action in an Friedmann–Robertson–Walker (FRW) universe constructed from products of any four-dimensional covariant tensors with free upper 0 indices, in the leading order, we can write out
@@
Neglecting the linear term above certain energy [1], now we construct the interactions when we couple the unparticles to the Goldstone $\pi$. In this work, we will be most interested in the interactions between the Goldstone and the scalar unparticles. We will also consider the couplings to the unparticle current and the stress tensor of unparticles. Since the unparticles are scaling invariant, the FRW curvature will not affect the conservation of the current, while it is natural to consider the traceless part of the stress tensor.
@@
where $a(t)$ is the FRW scale factor and we used the constraint $\nabla^{\mu}J_{\mu}=0$ at the background level, to replace $\dot{\pi}J_{0}$ by $J_{i}\partial_{i}\pi$. For spin-2 traceless stress tensor field $T_{\mu\nu}\equiv\mathcal{O}^{(2)}_{\mu\nu,\Delta=4}$, similar to the above discussion, the leading order interactions between $T_{\mu\nu}$ and the Goldstone $\pi$ will be
@@
where $f_{\pi}^{4}$ is the symmetry breaking scale [81] and we show the amplitude of scalar fluctuations that best fits the Planck satellite observations [82].
@@ 2.2 Strongly Coupled Sectors with Holographic Duals
It is of course difficult to treat strongly coupled field theories, as there are no general tools to study them. Besides going to the lattice, a very useful alternative is to use a weakly coupled dual description, if available. For some theories, the dual description is gravitational. We are considering a case in which we decouple gravity from the scalar dynamics during inflation, and consider a strongly coupled field theory in a rigid background. In that case, we can approximate the inflationary geometry to a fixed de Sitter background, and describe the holographic duals of strongly coupled quantum field theories in de Sitter space, see e.g. the discussion in [52, 83, 26].
@@
Now we consider mass deformations of the UV fixed point. The resulting gravity duals will depend on the dominant phase of the theory in the large $N$ limit [31, 36]. We will content ourselves with describing the gravitational duals that correspond to each phase, though at finite $N$ they will both be present and mixed by thermal fluctuations. The scenarios are similar to those described in [84, 68, 53, 83]. A useful setup is that of a five-dimensional CFT compactified on $\text{dS}_{4}\times S^{1}$. There are two saddles. The first one is the Schwarzschild geometry, which when Wick rotated gives a bubble of nothing spacetime [36, 40, 83], or a spacetime with an end of the world brane. This is the gravity dual of a gapped theory with mass much bigger than the Hubble scale of the boundary. As the mass gap is rather big, the states are inefficiently produced (being Boltzmann suppressed), so we expect the cosmological correlator in this phase to be given by tree-level exchange of the lightest state of the spectrum—which, being heavy, will be well-approximated by the equilateral shape of non-Gaussianity.
@@
Finally, when the mass deformation is not too large, there is a phase of the theory with a gap around the Hubble scale. The gravity dual has a holographic renormalization group (RG) flow that can be continued past the location of the putative IR brane, effectively “frying it.” [^4] Thanks to Raman Sundrum for explaining the appearance of the horizon through “brane frying.” In that case, the dual develops a cosmology behind this brane, a crunching cosmology with open FRW slices—see the discussion in [83]. A simple model of this gravity dual is $\text{AdS}_{6}$ with an identification. The identification makes the compactified circle shrink to zero size at the crunch.
@@
From [12, 13], we know that we can relate the correlation functions of massless inflatons $\phi$ to the correlation functions of conformally coupled scalars $\varphi$ through weight-shifting operators. Thus we consider an effective action of a conformally coupled $\varphi$ coupled to an unparticle field of dimension $\Delta$:
@@
In inflationary cosmology, since we are computing the equal-time correlator on the late-time boundary, we go to spatial momentum space. From [66, 85], we know that in $d$-dimensional Euclidean space,
@@
after normalizing this expression by $(2\pi)^{2}$. If we take $\Delta=1$, (3.11) will reduce to the well-known propagator of conformally coupled scalars [10]:
@@
In the “in-in” formalism (also called Schwinger-Keldysh formalism), the four-point correlation function with two-site tree-level exchange in the $s$-channel can always be divided into four sectors [86, 87, 88, 89, 11, 6, 10]:
@@ 3.1 Mellin-Barnes Integration
To make the notation consistent with [12], we map $X_{1}\mapsto 1/u$ and $X_{2}\mapsto 1/v$. Now the physical regions of the kinematic variables are $u\in(0,1]$, $v\in(0,1]$. Based on the derivation explained in detail in Appendix B, using Mellin-Barnes integration, we obtain the analytic form of the de Sitter four-point function, given by
@@
When $\Delta$ is an integer, there will be UV-divergences in (3.22). Physically, these divergences occurring at integer values of $\Delta$ can be interpreted as the divergences in $n$-loop “banana” diagrams involving multiple propagators. This is explained in detail in [90]; the essential idea is that in free field theory one can build integer-valued unparticle operators from taking operator products of free fields. This procedure requires proper renormalization, which is reflected in the divergences of the unparticle exchange diagram. Moreover, [85] showed that when $\Delta-2\in\mathds{N}^{0}$, we need to regularize the 2-point function, because the Euclidean 2-point function has no Fourier transform and requires regularization and renormalization, which leads to anomalies. The UV-divergences can be absorbed in local counterterms. For four-point functions, the counterterms we are able to add to the Lagrangian will be the contact interaction $\lambda\varphi^{4}$ and its higher-derivatives. As we will illustrate with the following examples, these counterterms are able to absorb the UV divergences. From [12], we know that contact interactions from irrelevant operators have shapes that can be calculated by repeated application of $\Delta_{u}$ on the simplest correlator coming from a $\varphi^{4}$ self-interaction, given by $uv/(u+v)$, with
@@
To perform the renormalization for integer $\Delta$’s, one has to expand the Appell $F_{1}$ functions. With the help of the package Diogenes developed in [91], we can find $\epsilon$-expansions for Appell $F_{1}$ functions expanded near integer values. The $\Delta=1$ case is well-known [10, 12], it is given by